### v1

# Note on shifted primes with large prime factors

Abstract.
For any $0<c<1$ let
$T_{c}(x)=|\big\{p\leq x:p\in\mathbb{P},P^{+}(p-1)\geq p^{c}\big\}|,$
where $\mathbb{P}$ is the set of primes and $P^{+}(n)$ denotes the largest prime factor of $n$. Erdős proved in 1935 that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\rightarrow 0,\quad\text{as~}c\rightarrow 1,$
where $\pi(x)$ denotes the number of primes not exceeding $x$. Recently, Ding gave a quantitative form of Erdős’ result and showed that for $8/9<c<1$ we have
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq 8\big(c^{-1}-1\big).$
In this article, Ding’s bound is improved to
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq-\frac{7}{2}\log c$
for $e^{-\frac{2}{7}}<c<1$.
Key words and phrases:
shifted primes, largest prime factor, linear sieve, primes in arithmetic progressions
2010 Mathematics Subject Classification:
Primary 11N05; Secondary 11N36
1. Introduction
We denote by $P^{+}(n)$ the largest prime factor of an integer $n$, with the convention that $P^{+}(1)=1$. The study of largest prime factor of shifted prime $P^{+}(p+a),a\in\mathbb{Z}$ is of significant importance. First, the infinitude of primes $p$ with $P^{+}(p+2)>p$ is equivalent to the twin prime conjecture, which is one of the most well-known open problems in number theory; second, an unexpected connection between large value of $P^{+}(p-1)$ and the first case of Fermat’s last theorem, was established by Adleman and Heath-Brown [1], and Fouvry [12]. Last but not least, small values of $P^{+}(p+a)$ plays an important role in cryptography, such as Pollard’s p1p-1 algorithm and Williams’ $p+1$ algorithm.
In this article, we study the quantity $T_{c}(x)$ defined by
$T_{c}(x):=\big|\big\{p\leq x:p\in\mathbb{P},P^{+}(p-1)\geq p^{c}\big\}\big|,$
where $0<c<1$ and $\mathbb{P}$ is the set of primes. As an application of the Bombieri-Vinogradov theorem as well as the Brun-Titchmarsh inequality, Goldfeld [13] proved in 1969 that
$\liminf_{x\rightarrow\infty}T_{1/2}(x)/\pi(x)\geq 1/2.$
Goldfeld also pointed out that his arguments could also lead to, for any $c<7/12$
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)>0.$ (1.1)
There are a number of improvements on the value of $c$ in (1.1), see, e.g. Motohashi [23], Hooley [20, 21], Deshouillers–Iwaniec [7], Fouvry [12] and Baker-Harman [2]. The best record of $c$ up to now is $0.677$, obtained by Baker and Harman [3].
In 2015 Luca, Menares and Pizarro-Madariaga [19] considered the explicit lower bound of $T_{c}(x)$ for small values of $c$. Specifically, for $1/4\leq c\leq 1/2$ they proved that
$T_{c}(x)\geq(1-c)\frac{x}{\log x}+E(x)$ (1.2)
where
$E(x)\ll\begin{cases}x\log\log x/(\log x)^{2},&\text{for~}1/4<c\leq 1/2,\\ x/(\log x)^{5/3},&\text{for~}c=1/4.\end{cases}$
Later, Chen and Chen [6] extended the range of $c$ to $(0,1/2)$ in (1.2) with slightly better $E(x)$. Chen and Chen also proved that for any $k\geq 2$ there exists at most one $c\in\big[\frac{1}{k+1},\frac{1}{k}\big)$ such that
$T_{c}(x)=(1-c)\frac{x}{\log x}+o\Big(\frac{x}{\log x}\Big).$
Based on their result, Chen and Chen conjectured that for any $k\geq 1$ and $c\in\big[\frac{1}{k+1},\frac{1}{k}\big)$ we have
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)\geq 1-\frac{1}{k+1}.$ (1.3)
In 2018, Feng and Wu [11] proved that
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)\geq 1-4\int_{1/c-1}^{1/c}\frac{\rho(t)}{t}dt$
holds for $0<c<0.3517\ldots$, where $\rho(u)$ is the Dickman function, defined as the unique continuous solution of the equation differential-difference
$\begin{cases}\rho(u)=1,&0\leq u\leq 1,\\ u\rho^{\prime}(u)=-\rho(u-1),&u>1.\end{cases}$
As a corollary, Feng and Wu proved conjecture (1.3) for $k\geq 3$ by numerical values involving the Dickman function. The lower bounds of $T_{c}(x)$ were further improved by Liu, Wu and Xi [18] to
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)\geq 1-4\rho(1/c)$
provided $0<c<0.3734\ldots$.
Later Ding [8, Final remarks] pointed out that conjecture (1.3) of Chen and Chen in fact contradicts the Elliott–Halberstam conjecture according to the works of Pomerance [24], Granville [14], Wang [25] and Wu [27]. That is, one has
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)=\lim_{x\rightarrow\infty}T_{c}(x)/\pi(x)=\left(1-\rho\left(\frac{1}{c}\right)\right)\rightarrow 0,\quad\text{as}~c\rightarrow 1$
under the assumption of the Elliott–Halberstam conjecture. Motivated by this, Ding [8] then proved unconditionally that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<1/2$ (1.4)
for some absolute constant $c<1$, thus disproving conjecture (1.3) for the case $k=1$. The proof of (1.4) by Ding is based on the following corollary of Brun-Titchmarsh type. For $(a,m)=1$ let $\pi(x;m,a)$ denotes the number of primes $p$ not exceeding $x$ such that $p\equiv a\pmod{m}$.
Proposition 1.1.
[27, Lemma 2.2] There exist two functions $K_{2}(\theta)>K_{1}(\theta)>0$, defined on the interval $(0,17/32)$ such that for each fixed $A>0$, and sufficiently large $Q=x^{\theta}$, the inequalities
$K_{1}(\theta)\frac{\pi(x)}{\varphi(m)}\leq\pi(x;m,1)\leq K_{2}(\theta)\frac{\pi(x)}{\varphi(m)}$
hold for all integers $m\in(Q,2Q]$ with at most $O\left(Q(\log Q)^{-A}\right)$ exceptions, where the implied constant depends only on $A$ and $\theta$. Moreover, for any fixed $\varepsilon>0$, these functions can be chosen to satisfy the following properties:
$\bullet$ $K_{1}(\theta)$ is monotonic decreasing, and $K_{2}(\theta)$ is monotonic increasing.
$\bullet$ $K_{1}(1/2)=1-\varepsilon$ and $K_{2}(1/2)=1+\varepsilon$.
The constant $c$ in (1.4) could further be specified by explicit values of $K_{1}(\theta)$ in Proposition 1.1. In fact, one has $K_{1}(\theta)\geq 0.16$ for $1/2\leq\theta\leq 13/25$ [2, Theorem 1] and $K_{1}(\theta)\geq 1/100$ for $13/25\leq\theta\leq 17/32$ [22, Eq. (4)]. Using the method of Ding [8] as well as the explicit values of $K_{1}(\theta)$, Xinyue Zang (private communication) obtained that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq\frac{0.496875}{c}<\frac{1}{2},\quad\text{for~}0.993375<c<1.$ (1.5)
However, there are earlier results related to conjecture (1.3) of Chen and Chen as well as Ding’s result (1.4). Actually, as indicated by the proof of a former result of Erdős [10, from line -6, page 212 to line 4, page 213], as early as 1935, people could already conclude that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\rightarrow 0,\quad\text{as}~c\rightarrow 1$ (1.6)
by combining with Wu’s lemma (see Lemma 2.3 below). Clearly, (1.4) is now a simple corollary of (1.6). In a later article, Ding [9] obtained a quantitative form of Erdős’ result (1.6), stating
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq 8\big(c^{-1}-1\big)$ (1.7)
for $8/9<c<1$. By (1.7) one notes easily that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<\frac{1}{2}$ (1.8)
for any $16/17<c<1$ which improved the numerical values of (1.5). It should be mentioned that almost the same time as Ding’s result (1.7), Bharadwaj and Rodgers [4] independently obtained the same result (1.6) with a general form in probabilistic language.All of the authors (Ding, Bharadwaj and Rodgers) were unaware of Erdős’ result at an earlier time. Erdős’ result (1.6) is an application of Brun’s method, while the proof of (1.7) is mainly based on the following quantitative version of Selberg’s upper bound sieve.
Proposition 1.2.
[15, page 172, Theorem 5.7] Let $g$ be a natural number, and let $a_{i},b_{i}~(i=1,2,\cdot\cdot\cdot,g)$ be integers satisfying
$E:=\prod_{i=1}^{g}a_{i}\prod_{1\leq r<s\leq g}(a_{r}b_{s}-a_{s}b_{r})\neq 0.$
Let $\varrho(p)$ denote the number of solutions $n\pmod{p}$ to the congruence
$\prod_{i=1}^{g}(a_{i}n+b_{i})\equiv 0\pmod{p}$,
and suppose that
$\varrho(p)<p\quad\text{for~all~}p.$
Let $y$ and $z$ be real numbers satisfying $1<y\leq z.$ Then we have
$\big|\big\{n:z-y<n\leq z,a_{i}n+b_{i}~\text{prime~for~}i=1,2,\cdot\cdot\cdot,g\big\}\big|\\ \leq 2^{g}g!\prod_{p}\left(1-\frac{\varrho(p)-1}{p-1}\right)\left(1-\frac{1}{p}\right)^{-g+1}\!\!\!\frac{y}{\log^{g}y}\left(1+O\left(\frac{\log\log 3y+\log\log 3|E|}{\log y}\right)\right),$
where the constant implied by the $O$-symbol depends at most on $g$.
For the proof of (1.7), one used Proposition 1.2 in the particular case $g=2$. Hence, the constant factor $8$ in (1.7) comes from the identity $2^{g}g!=8.$
In this article, we shall give a further improvement of (1.7) with two new ingredients: the first is the employment of Rosser-Iwaniec linear sieve to the prime variable sequence, instead of integer variable polynomial combining with the two dimensional sieve (i.e. Proposition 1.2 above); the second one is that when dealing with the error term coming from linear sieve, we apply a theorem of Bombieri-Friedlander-Iwaniec type with level of distribution $x^{4/7-\varepsilon}$ instead of the classical level $x^{1/2-\varepsilon}$.
Our main result is stated as follows.
Theorem 1.1.
For any $e^{-\frac{2}{7}}<c<1$ we have
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq-\frac{7}{2}\log c.$
Remark 1.
Theorem 1.1 provides a nontrivial upper bound of $T_{c}(x)$ for any $e^{-\frac{2}{7}}<c<1$. Here the lower bound of $c$ is approximately $e^{-\frac{2}{7}}=0.75147\ldots,$ which could be compared to $8/9=0.88888\ldots$ in (1.7). Thus, Theorem 1.1 extends the range of $c$ in (1.7). Furthermore, one may see easily that
$\frac{7}{2}\log c<8\big(c^{-1}-1\big)$
for any $8/9<c<1$ and hence Theorem 1.1 also improves the upper bound of $T_{c}(x)$ in (1.7).
The following corollary of Theorem 1.1 improved (1.4), (1.5) and (1.8) considerably.
Corollary 1.1.
For any $c>e^{-\frac{1}{7}}$ we have $\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<1/2$.
Remark 2.
The numerical value of $e^{-\frac{1}{7}}$ is $0.86687\ldots$. In [9, Remarks], it was concluded that $\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<1/2$ for any $c>e^{-\frac{1}{2}}=0.60653\ldots$ under the Elliott-Halberstam conjecture. Corollary 1.1 makes some further progress toward this direction.
2. Fundamental lemmas
Let $\mu(n)$ be the Möbius function. Let $\mathcal{A}$ be a finite sequence of positive integers and $\mathcal{P}$ a subset of primes. For any $z\geq 2$, let
$P(z)=\prod_{\begin{subarray}{c}p\leq z\\ p\in\mathcal{P}\end{subarray}}p.$
Next for square-free number $d$ with $d|P(z)$, we define
$\mathcal{A}_{d}=:\big\{a\in\mathcal{A}:d|a\big\}.$
Define the sieve function $S(\mathcal{A},\mathcal{P},z)$ to be
$S(\mathcal{A},\mathcal{P},z):=\big|\big\{a\in\mathcal{A}:(a,P(z))=1\big\}\big|.$
Suppose that $\big|\mathcal{A}_{d}\big|$ possesses the following form
$\big|\mathcal{A}_{d}\big|=\frac{\omega(d)}{d}X+r\big(\mathcal{A},d\big),$
where $X$ is an approximation to $|\mathcal{A}|$ and $\omega(d)$ is a multiplicative function satisfying
$0<\omega(p)<p,\quad p\in\mathcal{P}.$ (2.1)
Here, $\omega(d)d^{-1}X$ can be viewed as an approximation of the quantity $\big|\mathcal{A}_{d}\big|$ and $r\big(\mathcal{A},d\big)$ is regarded as the oscillation between $\big|\mathcal{A}_{d}\big|$ and $\omega(d)d^{-1}X$. We also let
$V(z)=:\prod_{\begin{subarray}{c}p|P(z)\end{subarray}}\left(1-\frac{\omega(p)}{p}\right).$
The first lemma is a result of Iwaniec [16, 17] on linear sieve with well factorable error terms. An arithmetic function $\lambda(q)$ is called well factorable of level $Q$ if for any $Q=Q_{1}Q_{2}$, $Q_{1},Q_{2}\geq 1$, there exist two functions $\lambda_{1}$ and $\lambda_{2}$ supported in $1,Q_{1}]$ and $1,Q_{2}]$ respectively such that
$|\lambda_{1}|\leq 1,\quad|\lambda_{2}|\leq 1\quad\text{and}\quad\lambda=\lambda_{1}\ast\lambda_{2}.$
Lemma 2.1.
Suppose that there is a constant $K\geq 2$ such that
$\prod_{\begin{subarray}{c}u\leq p<v\\ p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}\leq\frac{\log v}{\log u}\left(1+\frac{K}{\log u}\right)$
for any $v>u\geq 2$. Then for any $\varepsilon>0$ and $D^{1/2}\geq z\geq 2$ we have
$S(\mathcal{A},\mathcal{P},z)\leq XV(z)\left(F\left(\frac{\log D}{\log z}\right)+E\right)+\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}_{h}(d)r\big(\mathcal{A},d\big),$
where $sF(s)=2e^{\gamma}$ $(0<s\leq 3)$, $\gamma$ is the Euler constant, and the first error term $E$ satisfies
$E=O\big(\varepsilon+\varepsilon^{-8}e^{K}(\log D)^{-1/3}\big).$
The coefficients $\lambda^{+}_{h}(d)$ satisfy $\big|\lambda_{h}^{+}(d)\big|\leq 1$ and vanish for $d>D$ or $\mu(d)=0$. Especially, $\lambda^{+}_{h}(d)$ are well factorable of level $D$.
Let $\mathbb{N}$ be the set of natural numbers. For $q\in\mathbb{N}$ and $(a,q)=1$, define
$\pi(y;\ell,a,q)=\sum_{\begin{subarray}{c}\ell p\leq y\\ \ell p\equiv a\!\!\!\!\!\pmod{q}\end{subarray}}1,$
where the symbols $p$ will always be primes. As usual, let $\varphi(n)$ be the Euler totient function and ${\rm li}(y)=\int_{2}^{y}\frac{1}{\log t}dt$ be the Gauss function. The second lemma is the following theorem of Bombieri-Friedlander-Iwaniec type given by Wang [26, Proposition 3.2].
Lemma 2.2.
Let $a\neq 0$ be a given integer, and let $A>0$ and $\varepsilon>0$. For any well factorable function $\lambda(q)$ of level $Q$, the following estimate
$\sum_{\begin{subarray}{c}q\leq Q\\ (a,q)=1\end{subarray}}\lambda(q)\sum_{\begin{subarray}{c}L_{1}\leq\ell\leq L_{2}\\ (\ell,q)=1,~2|\ell\end{subarray}}\bigg(\pi(x;\ell,a,q)-\frac{{\rm li}\big(x/\ell\big)}{\varphi(q)}\bigg)\ll\frac{x}{(\log x)^{A}}$
holds for $Q\leq x^{4/7-\varepsilon}$ and $1\leq L_{1}\leq L_{2}\leq x^{1-\varepsilon}$, where the implied constants depend only on $a,A$ and $\varepsilon$.
Remark 3.
In fact, the original statement of Wang’s proposition [26, Proposition 3.2] is slightly different from Lemma 2.2. For our applications, we add the additional restriction $2|\ell$ here in Lemma 2.2. The proof is almost the same as Wang [26, Proposition 3.2]. After applying Heath-Brown’s identity, we shall consider the sum
$\Delta(L\mid M_{1},\ldots,M_{j}\mid N_{1},\ldots,N_{j};\,q,\,a)$
$\qquad:=\mathop{{\sum}^{*}}_{\begin{subarray}{c}2\ell m_{1}\ldots m_{j}n_{1}\ldots n_{j}\equiv a({\rm mod}\,q)\\ \ell\in\mathscr{L},\,m_{i}\in\mathscr{M}_{i},\,n_{i}\in\mathscr{N}_{i}\end{subarray}}\mu(m_{1})\ldots\mu(m_{j})-\frac{1}{\varphi(q)}\mathop{{\sum}^{*}}_{\begin{subarray}{c}(2\ell m_{1}\ldots m_{j}n_{1}\ldots n_{j},\,q)=1\\ \ell\in\mathscr{L},\,m_{i}\in\mathscr{M}_{i},\,n_{i}\in\mathscr{N}_{i}\end{subarray}}\mu(m_{1})\ldots\mu(m_{j}),$
where $\sum^{*}$ means that the summation is restricted to numbers $m_{1},\ldots,m_{j},n_{1},\ldots,n_{j}$ free of prime factors <z<z and $\mathscr{L},\mathscr{M}_{i},\mathscr{N}_{i}$ are intervals of the type
$\mathscr{L}:=[(1-\Delta)L,\,L[,\quad\mathscr{M}_{i}:=[(1-\Delta)M_{i},\,M_{i}[,\quad\mathscr{N}_{i}=[(1-\Delta)N_{i},\,N_{i}[$
with
$LM_{1}\ldots M_{j}N_{1}\ldots N_{j}=x,\qquad\max(M_{1},\ldots,M_{j})<x^{1/7}$
and $\Delta=(\log x)^{-A_{1}}$. Here $A_{1}$ is a sufficiently large constant. In the case $L=x^{\nu_{0}}\geqslant x^{3/7}$, we apply Theorem 5 of [5] with $M=L$, and here the coefficient $2$ is attached to $m_{1}\cdots m_{j}n_{1}\cdots n_{j}$. Otherwise, we shall apply Theorems 1, 2 and $5^{*}$ separately according to the partial product of $M_{1},\ldots,M_{j},N_{1},\ldots,N_{j}$ is located in some given intervals, and in these cases the coefficient $2$ is attach to $\ell$.
It seems that we may further generalize Lemma 2.2 to
$\sum_{(a,\,q)=1}\lambda(q)\sum_{\begin{subarray}{c}L_{1}\leqslant\ell\leqslant L_{2}\\ (\ell,\,q)=1\end{subarray}}f(\ell)\bigg(\pi(x;\,\ell,a,q)-\frac{{\rm{li}}(x/\ell)}{\varphi(q)}\bigg)\ll_{a,A,\varepsilon}\frac{x}{(\log x)^{A}}$
for some smooth function $f(\ell)\ll\tau(\ell)^{B}$ with $B>0$. The main difference between the proofs is that we need an analogue of Theorem 5 in [5] with coefficient $\alpha_{\ell}\equiv 1$ replaced by smooth function $f(\ell)$, which is just Bombieri-Friedlander-Iwaniec have done in the proof of Theorem 5 in [5]. Here we do not pursue the details.
The last lemma is another conjecture of Chen and Chen [6] which was later confirmed by Wu [27, Theorem 2].
Lemma 2.3.
For $0<c<1$, let
$T^{\prime}_{c}(x)=\#\big\{p\leq x:p\in\mathbb{P},P^{+}(p-1)\geq x^{c}\big\}.$
Then for sufficiently large $x$ we have
$T_{c}(x)=T^{\prime}_{c}(x)+O\left(\frac{x\log\log x}{(\log x)^{2}}\right).$
3. Proof of Theorem 1.1
Now, we turn to the proof of our theorem. Throughout, the symbols $p$ and $p^{\prime}$ will always be primes and $x$ is supposed to be sufficiently large.
First, by Lemma 2.3, it suffices to show that for any $e^{-\frac{2}{7}}<c<1$,
$\limsup_{x\rightarrow\infty}T^{\prime}_{c}(x)/\pi(x)\leq-\frac{7}{2}\log c.$
Clearly, for $c>e^{-\frac{2}{7}}>0.75$ we have
$T^{\prime}_{c}(x)=\sum_{\begin{subarray}{c}p^{\prime}\leq x\\ P^{+}(p^{\prime}-1)\geq x^{c}\end{subarray}}1=\sum_{x^{c}\leq p<x}\sum_{\begin{subarray}{c}p^{\prime}\leq x\\ p|p^{\prime}-1\end{subarray}}1=\sum_{x^{c}\leq p<x}\sum_{\begin{subarray}{c}\ell p+1\leq x\\ \ell p+1\in\mathcal{P}\\ 2|\ell\end{subarray}}1\leq\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\sum_{\begin{subarray}{c}\ell p\leq x\\ \ell p+1\in\mathcal{P}\end{subarray}}1.$
We are leading to sieve out primes in the following sequence
$\mathcal{A}:=\Big\{\ell p+1:\ell\leq x^{1-c},\ell p\leq x,2|\ell\Big\},$
where $c$ is a fixed number satisfying $0.75<c<1$. Let $\mathcal{P}=\mathbb{P}\setminus\{2\}$ and define the sieve function $S(\mathcal{A},\mathcal{P},z)$ to be
$S(\mathcal{A},\mathcal{P},z):=\big\{a\in\mathcal{A}:\big(a,P(z)\big)=1\big\},$
where $P(z)=\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}p$ and $z\leq x^{1/2}$ is a parameter to be decided later. Then, we deduce from above notation that
$T^{\prime}_{c}(x)\leq\sum_{\begin{subarray}{c}a\in\mathcal{A}\\ a\in\mathcal{P}\end{subarray}}1\leq S(\mathcal{A},\mathcal{P},z)+\pi(z)=S(\mathcal{A},\mathcal{P},z)+O\big(x^{1/2}\big).$ (3.1)
For applications of Lemma 2.1, we now need to specify $\big|\mathcal{A}_{d}\big|$ and $r\big(\mathcal{A},d\big)$. Let $\varepsilon>0$ be an arbitrary small number and $D=x^{4/7-\varepsilon}$. For any $d\leq D$ with $d|P(z)$, we have
$\big|\mathcal{A}_{d}\big|=\sum_{\begin{subarray}{c}\ell p\leq x\\ \ell\leq x^{1-c},\,2|\ell\\ \ell p+1\equiv 0\!\!\!\!\!\pmod{d}\end{subarray}}1=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\ \sum_{\begin{subarray}{c}\ell p\leq x\\ \ell p\equiv-1\!\!\!\!\!\pmod{d}\end{subarray}}1=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\pi(x;\ell,-1,d).$
Next, we naturally approximate $\pi(x;\ell,-1,d)$ by ${\rm li}(x/\ell)/\varphi(d)$ and we write
$\big|\mathcal{A}_{d}\big|=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}+r_{1}(\mathcal{A},d),$ (3.2)
where $r_{1}(\mathcal{A},d)$ is the error term:
$r_{1}(\mathcal{A},d)=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{{\rm li}(x/\ell)}{\varphi(d)}\bigg).$ (3.3)
Now we turn to estimate the sum over $\ell$ in (3.2), where the main term comes from.
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}$ $=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{1}{\varphi(d)}\int_{2}^{x/\ell}\frac{{\rm{d}}t}{\log t}$
$=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{x/\ell}{\varphi(d)\log(x/\ell)}\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}$
$=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{\varphi(d)}\sum_{\begin{subarray}{c}\ell\leq x^{1-c}/2\\ (2\ell,\,d)=1\end{subarray}}\ \frac{1}{2\ell\log(x/\ell)}.$
The condition (2,d)=1(2,\,d)=1 is in fact redundant since $d|P(z)=\prod_{2<p\leq z}p$. To relax the condition $(\ell,\,d)=1$, we employ the Möbius inversion getting
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}$ $=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{2\varphi(d)}\sum_{\begin{subarray}{c}\ell\leq x^{1-c}/2\end{subarray}}\ \frac{1}{\ell\log(x/\ell)}\sum_{e\mid(\ell,\,d)}\mu(e)$
$=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{2\varphi(d)}\sum_{e\mid d}\,\frac{\mu(e)}{e}\sum_{e\ell<x^{1-c}/2}\frac{1}{\ell\log(x/e\ell)}$
$=\Big\{1+O\Big(\frac{1}{\log x}\Big)\Big\}\frac{x}{2\varphi(d)}\big(S_{1}+S_{2}\big),$ (3.4)
where
$S_{1}:=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\sum_{\ell<x^{1-c}/2e}\frac{1}{\ell\log(x/e\ell)},\qquad S_{2}:=\sum_{\begin{subarray}{c}e\mid d\\ e\geq(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\sum_{\ell<x^{1-c}/2e}\frac{1}{\ell\log(x/e\ell)}.$
First, we estimate $S_{2}$ trivially
$S_{2}\ll\sum_{\begin{subarray}{c}e\mid d\\ e\geq(\log x)^{9}\end{subarray}}\frac{1}{e}\ll\frac{\tau(d)}{(\log x)^{9}},$ (3.5)
where $\tau(d)$ denotes the number of divisors of $d$.
And for $S_{1}$, we have by the partial summation
$S_{1}$ $=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg\{\int^{x^{1-c}/2e}_{1}\frac{{\rm{d}}t}{t\log(x/et)}+O\Big(\frac{1}{\log x}\Big)\bigg\}$
$=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg\{\log\bigg(\frac{\log(x/e)}{c\log x}\bigg)+O\Big(\frac{1}{\log x}\Big)\bigg\}$
$=\sum_{\begin{subarray}{c}e\mid d\\ e<(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg\{\log\frac{1}{c}+O\bigg(\frac{\log\log x}{\log x}\bigg)\bigg\},$
where we have removed $\log e$ with an admissible error term in the last step thanks to the condition $e<(\log x)^{9}$. Now we reinsert the sum over $e$ with $e\geq(\log x)^{9}$ up to an error term as in (3.5) getting
$S_{1}$ $=\bigg(\sum_{e\mid d}\,\frac{\mu(e)}{e}-\sum_{\begin{subarray}{c}e\mid d\\ e\geq(\log x)^{9}\end{subarray}}\frac{\mu(e)}{e}\bigg)\bigg\{\log\frac{1}{c}+O\bigg(\frac{\log\log x}{\log x}\bigg)\bigg\}$
$=\frac{\varphi(d)}{d}\bigg\{\log\frac{1}{c}+O\bigg(\frac{\log\log x}{\log x}\bigg)\bigg\}+O\bigg(\frac{\tau(d)}{(\log x)^{9}}\bigg).$ (3.6)
Combining (3), (3.5) and (3) we obtain
$\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\frac{{\rm li}(x/\ell)}{\varphi(d)}=\frac{-\log c}{2}x\cdot\frac{1}{d}\Big\{1+O\Big(\frac{\log\log x}{\log x}\Big)\Big\}+O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg),$
whence by (3.2) and (3.3) we arrive at
$|\mathcal{A}_{d}|=\frac{\omega(d)}{d}X+r(\mathcal{A},d)$
where
$X$ $=\frac{-\log c}{2}x\Big\{1+O\Big(\frac{\log\log x}{\log x}\Big)\Big\},\quad\omega(d)=1,$
and
$r(\mathcal{A},d)$ $=r_{1}(\mathcal{A},d)+O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg)$
$=\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,\,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{{\rm li}(x/\ell)}{\varphi(d)}\bigg)+O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg).$
Now we are ready to apply Lemma 2.1. First we need to verify that the condition
$\prod_{\begin{subarray}{c}u\leq p<v\\ p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}\leq\frac{\log v}{\log u}\left(1+\frac{K}{\log u}\right)$ (3.7)
holds for some absolute constant $K$. In fact, by Mertens’ formula it is easy to see
$\prod_{\begin{subarray}{c}u\leq p<v\\ p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}=\prod_{\begin{subarray}{c}u\leq p<v\\ p>2\end{subarray}}\left(1-\frac{1}{p}\right)^{-1}=\frac{\log v}{\log u}\bigg(1+O\Big(\frac{1}{\log u}\Big)\bigg).$
Hence, with the choices of $X,\omega(d),r(\mathcal{A},d)$ as above, we deduce from Lemma 2.1 by taking $D=x^{4/7-\varepsilon}$ and $z=D^{\frac{1}{2}}=x^{\frac{2}{7}-\frac{\varepsilon}{2}}$ that
$S(\mathcal{A},\mathcal{P},z)$ $\leq XV(z)\bigg(F\Big(\frac{\log D}{\log z}\Big)+E\bigg)+\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}_{h}(d)r\big(\mathcal{A},d\big),$
$=\frac{-\log c}{2}x\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(1-\frac{1}{p}\Big)\Big(F(2)+o(1)\Big)+\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}_{h}(d)r_{1}\big(\mathcal{A},d\big)$
$\qquad+\sum_{h<\exp(8/\varepsilon^{2})}\,\sum_{d|P(z),\,d\leq D}O\bigg(\frac{x\tau(d)}{\varphi(d)(\log x)^{9}}\bigg)$ (3.9)
$=:S_{M}+S_{E1}+S_{E2},$
say. For the main term $S_{M}$, employing again Mertens’ formula
$\prod_{\begin{subarray}{c}2<p\leq z\end{subarray}}\Big(1-\frac{1}{p}\Big)=\prod_{\begin{subarray}{c}2<p\leq x^{\frac{2}{7}-\frac{\varepsilon}{2}}\end{subarray}}\Big(1-\frac{1}{p}\Big)=\frac{2e^{-\gamma}+o(1)}{\log(x^{2/7-\varepsilon/2})}=\Big(7e^{-\gamma}+o(1)\Big)\frac{1}{\log x},$
and noting that $F(2)=e^{\gamma}$, we obtain
$S_{M}$ $=\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x}.$ (3.10)
We are now in a position to apply Lemma 2.2 to estimate the first error term $S_{E1}$, provided the function $\lambda^{+}_{h}(d)$ is well factorable of level $D=x^{4/7-\varepsilon}$. By taking $L_{1}=1$ and $L_{2}=x^{1-c}$ in Lemma 2.2, we have
$S_{E1}$ $=\sum_{h<\exp(8/\varepsilon^{2})}\sum_{\begin{subarray}{c}d\leq x^{4/7-\varepsilon}\end{subarray}}\lambda^{+}_{h}(d)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,(\ell,d)=1\end{subarray}}r_{1}\big(\mathcal{A},d\big)$
$=\sum_{h<\exp(8/\varepsilon^{2})}\sum_{\begin{subarray}{c}d\leq x^{4/7-\varepsilon}\end{subarray}}\lambda^{+}_{h}(d)\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell,(\ell,d)=1\end{subarray}}\bigg(\pi(x;\ell,-1,d)-\frac{{\rm li}\big(x/\ell\big)}{\varphi(d)}\bigg)$
$\ll_{\varepsilon}\frac{x}{(\log x)^{A}}$ (3.11)
for any $A>0$, which is admissible.
For the error term $S_{E2}$, it is easy to see
$S_{E2}\ll\frac{x}{(\log x)^{9}}\sum_{d\leq x^{4/7-\varepsilon}}\frac{\tau(d)}{\varphi(d)}\ll\frac{x}{(\log x)^{6}},$ (3.12)
which is also admissible.
Then, inserting (3.10), (3) and (3.12) into (LABEL:SM+SE1+SE2), we arrive at
$S(\mathcal{A},\mathcal{P},z)\leq\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x}.$
Finally, $T^{\prime}_{c}(x)$ is estimated from (3.1) that
$T^{\prime}_{c}(x)\leq S(\mathcal{A},\mathcal{P},z)+O(x^{1/2})\leq\Big(-\frac{7}{2}\log c+o(1)\Big)\frac{x}{\log x},$
whence
$\limsup_{x\rightarrow\infty}T^{\prime}_{c}(x)/\pi(x)\leq-\frac{7}{2}\log c$
for any $e^{-\frac{2}{7}}<c<1$. This completes the proof of Theorem 1.1.
Acknowledgments
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFA1000700).
Y. C. was supported by National Natural Science Foundation of China (Grant No. 12201544) and China Postdoctoral Science Foundation (Grant No. 2022M710121).
Z. W. was supported by National Natural Science Foundation of China (Grant No. 12322101) and National Natural Science Foundation of Shandong Province (Grant No. ZR2023YQ004).